OpenStudy (anonymous):

-5(3-q)+4=5q-11 (PLEASE SHOW WORK)

jimthompson5910 (jim_thompson5910):

-5(3-q)+4=5q-11 -15+5q+4=5q-11 5q-11=5q-11 5q=5q-11+11 5q-5q=-11+11 0q=-11+11 0q=0 0=0 Since this equation is always true for any q value, this means q can equal any number. So there are an infinite number of solutions.